Bovar Company began the manufacture of new paging machines. The company installed a standard costing system to account for manufacturing costs. The standard cost per unit is:

              Direct materials   3 pound @ $5 per pound

              Direct labor         .5 hours @ $20 per direct labor hour

              Variable overhead       75% of direct labor cost

Actual production was 4,000 units; actual sales were 2,500 units.

There is no beginning inventory for direct materials.

Other data are:

  1. Materials price variance is $3,250 unfavorable.
  2. Materials efficiency variance is $2,500 favorable.
  3. Direct labor rate variance is $2,100 favorable.
  4. Direct labor efficiency variance is $2,000 unfavorable.
  5. Revenue was $125,000.
  6. Purchases of direct materials is $68,250.
  7. Actual variable overhead is $29,000.

REQUIRED:

Determine the following factors:

  1. actual quantity of materials used AND actual quantity of direct materials purchased
  2. actual direct materials price per pound
  3. variable overhead spending variance
  4. variable overhead efficiency variance
Answer #1

a)Material Price variance = Actual cost of quantity purchased - [AQ purchased *standard cost per pound]

        3250 = 68250 - [AQ purchased *5 ]

       AQ purchased * 5 = 68250 -3250

      AQ purchased = 65000/5

                              = 13000 pounds

Material efficiency variance = Standard cost per pound [AQ used- standard quantity for actual production]

    -2500 = 5 [AQ used - (4000*3)]

    -2500/5 = [AQ used - 12000]

     -500 = [AQ used -12000]

    AQ used = -500 +12000

                  = 11500 pounds

2)Actual direct material price per pound = Actual cost of material purchased / Actual quantity purchased

                    = 68250 /13000

                    = $ 5.25 per pound

3)Variable overhead spending variance = Actual variable overhead -[Actual hours *standard variable overhead rate per hour ]

               = 29000 - [2100 * 15 ]

              = 29000 - 31500

             = - 2500 F

**Working :

Direct Labor efficiency variance = Standard rate per hour [AH -Standard hours for actual production]

   2000 = 20 [AH -(4000*.5)]

2000/20 = [AH - 2000]

    100 +2000 =AH

Actual hours = 2100 hours

STANDARD OVERHEAD RATE = 20*75% = 15 PER HOUR

d)

Variable overhead efficiency variance =Standard rate per hour [Actual hours **standard hours for actual output]

                  = 15 [2100 - (4000*.50)]

                  =15[2100-2000]

                   = 15 *100

                    =1500 U